### v1

@@ Abstract
{IEEEkeywords}
Deep neural networks, Robustness, Signal detection, Unmanned aerial vehicles
1 INTRODUCTION
[…]
@@ 1.1 RELATED WORK
With the advent of DL-based methods, the data used to train models became the cornerstone of any detection system. Table 1 provides an overview of openly available datasets of RF drone signals. The DroneRF dataset [3] is one of the first openly available datasets. It contains RF time series data from three drones in four flight modes (i.e. on, hovering, flying, video recording) recorded by two universal software radio peripheral (USRP) SDR transceivers [4]. The dataset is widely used and enabled follow-up work with different approaches to classification systems, i.e. DL-based [5, 6], focused on pre-processing and combining signals from two frequency bands [7], genetic algorithm-based heterogeneous integrated k-nearest neighbour [8], and hierarchical reinforcement learning-based [9]. In general, the classification accuracies reported in the papers on the DroneRF dataset are close to $100\%$. Specifically, [4], [5], and [6] report an average accuracy of $99.7\%$, $100\%$, and $99.98\%$, respectively, to detect the presence of a drone. There is therefore an obvious need for a harder, more realistic dataset.
@@
Table 1: Overview on openly available drone RF datasets.
@@
1.2 MOTIVATION
As we have seen in other fields, such as computer vision, the success of DL can be attributed to: (a) high-capacity models; (b) increased computational power; and (c) the availability of large amounts of labelled data [22]. Thus, given the large amount of available raw RF signals (cf. Tab. 1) we promote the idea of open and reusable data, to facilitate model development and model comparison.
@@ 1.2 MOTIVATION
In the next section, we briefly describe the data collection and preprocessing procedure. Section 3 describes the model architectures and their training/validation method. In addition, we describe the setup of a low-cost drone detection system and of the field test. The resulting performance metrics are presented in Section 4 and are further discussed in Section 5.
2 MATERIALS
@@
2.1 DATA ACQUISITION
The drone’s remote control and, if present, the drone itself were placed in an anechoic chamber to record the raw RF signal without interference for at least one minute. The signals were received by a log-periodic antenna and sampled and stored by an Ettus Research USRP B210, see Fig. 1. In the static measurement, the respective signals of the remote control (TX) alone or with the drone (RX) were measured. In the dynamic measurement, one person at a time was inside the anechoic chamber and operated the remote control (TX) to generate a signal that is as close to reality as possible. All signals were recorded at a sampling frequency of $56$ MHz (highest possible real-time bandwidth). All drone models and recording parameters are listed in Tab. 2, including both uplink and downlink signals.
@@ 2.1 DATA ACQUISITION
Table 2: Transmitters and receivers recorded in the development dataset and their respective class labels. Additionally, we show the center frequency (GHz), the channel spacing (MHz), the burst duration (ms), and the repetition period of the respective signals (ms).
@@ 2.2 DATA PREPARATION
The drone signals occur in short bursts with some low power gain or background noise in between (cf. Tab. 2). We divided the signals into non-overlapping vectors of $1048576$ samples ($74.9$ ms) and only vectors containing a burst, or at least a partial burst, were used for the development dataset. This was achieved by applying an energy threshold. As the recordings were made in an echo-free chamber, the signal burst is always clearly visible. Hence, we only used vectors that contained a portion of the signal whose energy was above the threshold, which was arbitrarily set at $0.001$ of the average energy of the entire recording.
@@ 2.3 DEVELOPMENT DATASET
As described in Sec. 2.2, the drone signals were mixed with noise. More specifically, $50\%$ of the drone signals were mixed with Labnoise (Bluetooth + Wi-Fi + Amplifier) and $50\%$ with Gaussian noise. In addition, we created a separate noise class by mixing Labnoise and Gaussian noise in all possible combinations (i.e., Labnoise $+$ Labnoise, Labnoise $+$ Gaussian noise, Gaussian noise $+$ Labnoise, and Gaussian noise $+$ Gaussian noise). For the drone signal classes, as for the noise class, the number of samples for each SNR level was evenly distributed over the interval of SNRs $\in[-20,30]$ dB in steps of $2$ dB, i.e., $679$-$685$ samples per SNR level. The resulting number of samples per class is given in Tab. 3.
Table 3: Number of samples in the different classes in the development dataset.
@@
2.4 DETECTION SYSTEM PROTOTYPE
@@
3 METHODS
3.1 MODEL ARCHITECTURE AND TRAINING
@@ 3.2 MODEL EVALUATION
As accuracy simply measures the proportion of correct predictions out of the total number of observations, it can be misleading for unbalanced datasets. In our case, the noise class is over-represented in the dataset (cf. Tab. 3). Therefor, we also report the balanced accuracy, which is defined as the average of the recall obtained for each class, i.e. it gives equal weight to each class regardless of how frequent or rare it is.
3.3 VISUALISATION OF MODEL EMBEDDINGS
@@
3.4 DETECTION SYSTEM FIELD TEST
@@
Table 4 lists the drones and/or remote controls used in the field test. Note that the Graupner drone and remote control are part of the development dataset (cf. Tab. 2), but were not measured in the field experiment. We assume that no other drones were present during the measurements, so recordings where none of our transmitters were used are labelled as “Noise”.
Table 4: Drones and/or remotes used in the field test
@@
For each transmitter, distance, and angle, $20$ to $30$ s, or approximately $300$ spectrograms were live classified and recorded. The resulting number of samples for each class, distance, and angle are shown in Tab. 5.
Table 5: Number of samples (#samples) for each class, distance and antenna direction (angle) recorded in the field test. Recordings at 00 m distance have no active transmitter and were therefore labelled “Noise”
@@
4 RESULTS
4.1 CLASSIFICATION PERFORMANCE ON THE DEVELOPMENT DATASET
Table 6 shows the general mean $\pm$ standard deviation of accuracy and balanced accuracy on the test data of the development dataset (cf. Sec. 2.3), obtained in the $5$-fold cross-validation of the different models.
@@ 4.1 CLASSIFICATION PERFORMANCE ON THE DEVELOPMENT DATASET
Table 6: Mean $\pm$ standard deviation of the accuracy (Acc.) and the balanced accuracy (balanced Acc.) obtained in $5$-fold cross-validation of the different models on the test data of the development dataset. An indication of the model training time is given with the mean $\pm$ standard deviation of the number of training epochs (#epochs), i.e. when the highest balanced accuracy on the validation set was reached. The number of trainable parameters (#params) indicates the complexity of the model.
@@
4.2 EMBEDDING SPACE VISUALISATION
@@
4.3 CLASSIFICATION PERFORMANCE IN THE FIELD TEST
For each model architecture, we performed $5$-fold cross-validation on the development dataset (cf. Sec. 3.1), resulting in five trained models per architecture. Thus, we also evaluated all five trained models on the field test data. We report the balanced accuracy $\pm$ standard deviation for each model architecture for the complete field test dataset averaged over all directions and distances in Tab. LABEL:tab:field_test_acc.
Table 7: Mean $\pm$ standard deviation of the balanced accuracy (balanced Acc.) of the complete field test recordings for the different models.
@@ 4.3 CLASSIFICATION PERFORMANCE IN THE FIELD TEST
As observed on the development dataset (cf. Tab. 6), there is no meaningful difference in performance between the model architectures. We therefore focus on VGG11_BN, the simplest model trained, in the more detailed analysis of the field test results.
@@
Table 8: Mean balanced accuracy $\pm$ standard deviation of the VGG11_BN models on the field test recordings for the different classes for each direction ($0^{\circ}$, $90^{\circ}$ and $180^{\circ}$). The upper part shows the accuracies for the classification problem (seven classes) and the lower part the accuracies for the detection problem “Drone” or “Noise”
@@
Table 9: Mean balanced accuracy $\pm$ standard deviation of the VGG11_BN models on the field test data with active transmitters collected at different distances for each antenna direction ($0^{\circ}$, $90^{\circ}$ and $180^{\circ}$). The upper part shows the accuracies for the classification problem (seven classes) and the lower part the accuracies for the detection problem “Drone” or “Noise”
@@ 5 DISCUSSION
Obviously, the VGG models were able to learn the relevant features for the drone classification from the complex spectrograms of the RF signal. In this respect, we did not find any advantage for the use of more complex models, such as VGG19_BN, over the least complex model, VGG11_BN (cf. Tabs. 6 and LABEL:tab:field_test_acc).
@@
Another issue to consider is the occurrence of unknown drones, i.e. drones that are not part of the train set. Examining the embedding space (cf. 4.2) gives a first idea of whether a signal is clearly part of a known dense drone cluster or rather falls into the larger, less dense, noise cluster. We believe that a combination of an unsupervised deep autoencoder approach [33, 34] with an additional classification part (cf. [35]) would allow, first, to provide a stable classification of known samples and, second, to indicate whether a sample is known or rather an anomaly.

### v2

@@ Abstract
Index Terms:
Drone detection, UAV classification, Low signal-to-noise ratio, Robustness, Real-world field test.
I Introduction
@@ I-A Related Work
With the advent of DL-based methods, the data used to train models became the cornerstone of any detection system. Table I provides an overview of openly available datasets of RF drone signals. The DroneRF dataset [3] is one of the first openly available datasets. It contains RF time series data from three drones in four flight modes (i.e. on, hovering, flying, video recording) recorded by two universal software radio peripheral (USRP) SDR transceivers [4]. The dataset is widely used and enabled follow-up work with different approaches to classification systems, i.e. DL-based [5, 6], focused on pre-processing and combining signals from two frequency bands [7], genetic algorithm-based heterogeneous integrated k-nearest neighbour [8], and hierarchical reinforcement learning-based [9]. In general, the classification accuracies reported in the papers on the DroneRF dataset are close to $100\%$. Specifically, [4], [5], and [6] report an average accuracy of $99.7\%$, $100\%$, and $99.98\%$, respectively, to detect the presence of a drone. There is therefore an obvious need for a harder, more realistic dataset.
@@
TABLE I: Overview on openly available drone RF datasets.
@@
I-B Motivation
As we have seen in other fields, such as computer vision, the success of DL can be attributed to: (a) high-capacity models; (b) increased computational power; and (c) the availability of large amounts of labelled data [22]. Thus, given the large amount of available raw RF signals (cf. Tab. I) we promote the idea of open and reusable data, to facilitate model development and model comparison.
@@ I-B Motivation
In the next section, we briefly describe the data collection and preprocessing procedure. Section III describes the model architectures and their training/validation method. In addition, we describe the setup of a low-cost drone detection system and of the field test. The resulting performance metrics are presented in Section IV and are further discussed in Section V.
II Materials
@@
II-A Data Acquisition
The drone’s remote control and, if present, the drone itself were placed in an anechoic chamber to record the raw RF signal without interference for at least one minute. The signals were received by a log-periodic antenna and sampled and stored by an Ettus Research USRP B210, see Fig. 1. In the static measurement, the respective signals of the remote control (TX) alone or with the drone (RX) were measured. In the dynamic measurement, one person at a time was inside the anechoic chamber and operated the remote control (TX) to generate a signal that is as close to reality as possible. All signals were recorded at a sampling frequency of $56$ MHz (highest possible real-time bandwidth). All drone models and recording parameters are listed in Tab. II, including both uplink and downlink signals.
@@ II-A Data Acquisition
TABLE II: Transmitters and receivers recorded in the development dataset and their respective class labels. Additionally, we show the center frequency (GHz), the channel spacing (MHz), the burst duration (ms), and the repetition period of the respective signals (ms).
@@ II-B Data Preparation
The drone signals occur in short bursts with some low power gain or background noise in between (cf. Tab. II). We divided the signals into non-overlapping vectors of $1048576$ samples ($74.9$ ms) and only vectors containing a burst, or at least a partial burst, were used for the development dataset. This was achieved by applying an energy threshold. As the recordings were made in an echo-free chamber, the signal burst is always clearly visible. Hence, we only used vectors that contained a portion of the signal whose energy was above the threshold, which was arbitrarily set at $0.001$ of the average energy of the entire recording.
@@ II-C Development Dataset
As described in Sec. II-B, the drone signals were mixed with noise. More specifically, $50\%$ of the drone signals were mixed with Labnoise (Bluetooth + Wi-Fi + Amplifier) and $50\%$ with Gaussian noise. In addition, we created a separate noise class by mixing Labnoise and Gaussian noise in all possible combinations (i.e., Labnoise $+$ Labnoise, Labnoise $+$ Gaussian noise, Gaussian noise $+$ Labnoise, and Gaussian noise $+$ Gaussian noise). For the drone signal classes, as for the noise class, the number of samples for each SNR level was evenly distributed over the interval of SNRs $\in[-20,30]$ dB in steps of $2$ dB, i.e., $679$-$685$ samples per SNR level. The resulting number of samples per class is given in Tab. III.
TABLE III: Number of samples in the different classes in the development dataset.
@@
II-D Detection System Prototype
@@
III Methods
III-A Model Architecture and Training
@@ III-B Model Evaluation
As accuracy simply measures the proportion of correct predictions out of the total number of observations, it can be misleading for unbalanced datasets. In our case, the noise class is over-represented in the dataset (cf. Tab. III). Therefor, we also report the balanced accuracy, which is defined as the average of the recall obtained for each class, i.e. it gives equal weight to each class regardless of how frequent or rare it is.
III-C Visualisation of Model Embeddings
@@
III-D Detection System Field Test
@@
Table IV lists the drones and/or remote controls used in the field test. Note that the Graupner drone and remote control are part of the development dataset (cf. Tab. II), but were not measured in the field experiment. We assume that no other drones were present during the measurements, so recordings where none of our transmitters were used are labelled as “Noise”.
TABLE IV: Drones and/or remotes used in the field test
@@
For each transmitter, distance, and angle, $20$ to $30$ s, or approximately $300$ spectrograms were live classified and recorded. The resulting number of samples for each class, distance, and angle are shown in Tab. V.
TABLE V: Number of samples (#samples) for each class, distance and antenna direction (angle) recorded in the field test. Recordings at 00 m distance have no active transmitter and were therefore labelled “Noise”
@@
IV Results
IV-A Classification performance in the Development Dataset
Table VI shows the general mean $\pm$ standard deviation of accuracy and balanced accuracy on the test data of the development dataset (cf. Sec. II-C), obtained in the $5$-fold cross-validation of the different models.
@@ IV-A Classification performance in the Development Dataset
TABLE VI: Mean $\pm$ standard deviation of the accuracy (Acc.) and the balanced accuracy (balanced Acc.) obtained in $5$-fold cross-validation of the different models on the test data of the development dataset. An indication of the model training time is given with the mean $\pm$ standard deviation of the number of training epochs (#epochs), i.e. when the highest balanced accuracy on the validation set was reached. The number of trainable parameters (#params) indicates the complexity of the model.
@@
IV-B Embedding Space Visualisation
@@
IV-C Classification performance in the Field Test
For each model architecture, we performed $5$-fold cross-validation on the development dataset (cf. Sec. III-A), resulting in five trained models per architecture. Thus, we also evaluated all five trained models on the field test data. We report the balanced accuracy $\pm$ standard deviation for each model architecture for the complete field test dataset averaged over all directions and distances in Tab. LABEL:tab:field_test_acc.
TABLE VII: Mean $\pm$ standard deviation of the balanced accuracy (balanced Acc.) of the complete field test recordings for the different models.
@@ IV-C Classification performance in the Field Test
As observed on the development dataset (cf. Tab. VI), there is no meaningful difference in performance between the model architectures. We therefore focus on VGG11_BN, the simplest model trained, in the more detailed analysis of the field test results.
@@
TABLE VIII: Mean balanced accuracy $\pm$ standard deviation of the VGG11_BN models on the field test recordings for the different classes for each direction ($0^{\circ}$, $90^{\circ}$ and $180^{\circ}$). The upper part shows the accuracies for the classification problem (seven classes) and the lower part the accuracies for the detection problem “Drone” or “Noise”
@@
TABLE IX: Mean balanced accuracy $\pm$ standard deviation of the VGG11_BN models on the field test data with active transmitters collected at different distances for each antenna direction ($0^{\circ}$, $90^{\circ}$ and $180^{\circ}$). The upper part shows the accuracies for the classification problem (seven classes) and the lower part the accuracies for the detection problem “Drone” or “Noise”
@@ V Discussion
Obviously, the VGG models were able to learn the relevant features for the drone classification from the complex spectrograms of the RF signal. In this respect, we did not find any advantage for the use of more complex models, such as VGG19_BN, over the least complex model, VGG11_BN (cf. Tabs. VI and LABEL:tab:field_test_acc).
@@
Another issue to consider is the occurrence of unknown drones, i.e. drones that are not part of the train set. Examining the embedding space (cf. IV-B) gives a first idea of whether a signal is clearly part of a known dense drone cluster or rather falls into the larger, less dense, noise cluster. We believe that a combination of an unsupervised deep autoencoder approach [33, 34] with an additional classification part (cf. [35]) would allow, first, to provide a stable classification of known samples and, second, to indicate whether a sample is known or rather an anomaly.